factorize x 6 -10x 3 +27

factorize


x6-10x3+27

Grade:10

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

f(x) = x^6 - 10x^3 + 27
f(x) = x^3(x^3 - 10) + 27
f(x) = (x^2 + 2x + 3)(x^4 - 2x^3 + x^2 - 6x + 9)
All the roots are complex.
If you want you can factorise it furthur.